The contents of a certain box consist of 14 apples and 23 oranges. How many oranges must be removed from the box so that 70 percent of the pieces of fruit in the box will be apples?
This is what I have so far: 14 apples + 23 oranges = 37 total pieces of fruit; 14 apples/37-X=.70; then I get lost after this...

Answer by checkley75(3666)

14/(37-X)=.70 CROSS MULTIPLY
.7(37-X)=14
25.9-.7X=14
-.7X=14-25.9
-.7X=-11.9
X=-11.9/-.7
X=17 NUMBER OF ORANGES THAT SHOULD BE REMIOVED TO HAVE 70% OF THE FRUIT TO BE APPLES.
PROOF
14/(37-17)=.7
14/20=.7
.7=.7